1/1

Frédéric Bérat ,

My Normandy

tradukita de Tony Provencher

When far from us the Winter flies, When the world is born to hope anew, Under France’s lovely skies, When the sun returns in sweeter hue, When Nature ’round us greener be, When swallows homeward wing their way, I love to see my Normandy, The land that gave to me the light of day.

Switzerland’s dales I did behold, And her chalets and glaciers blue; I’ve seen Italian skies of gold, And the gondoliers of Venice too. Although these lands enchanted me, I knew there was no hideaway More lovely than my Normandy, The land that gave to me the light of day.

There comes a time, as life unfolds, That brings an end to reveries. A time when must the ransomed soul Revisit cherished memories. When time has chilled my muse and me, And songs of love are sung away, I’ll see again my Normandy, The land that gave to me the light of day.

Traduko de la Franca poemo “Ma Normandie” de Frédéric Bérat en la Anglan de Tony Provencher.

Arg-1029-2088 (2014-05-29 19:58:50)

La kanto estas aŭdebla en http: // manormandie. snoupi. com/ music . The website of Mr. Provencher can be found at http: // tonysmusic. snoupi. com/ home2 .

http://www.poezio.net/version?poem-id=1029&version-id=,2088